OpenStudy (anonymous):

Kenny bought 4 tickets to the baseball game. Each ticket cost $3. He gave the cashier $20. How much change would Kenny get? A. Calculate the amount of money that 4 tickets cost. Multiply that number by the amount of money Kenny gave the cashier. B. Calculate how much one ticket costs. Multiply that number by the amount of money Kenny gave the cashier. C. Calculate the amount of money that 4 tickets cost. Divide that number by the amount of money Kenny gave the cashier. D. Calculate the amount of money that 4 ticke

OpenStudy (anonymous):

4*3=12 12-20 dollars =8

OpenStudy (anonymous):

A. 12 x 20 = 240 B. 3 x 20 = 60 C. 12 / 20 = 0.6 D. 12 These directions make no sense.

OpenStudy (mrnood):

@gloriasanchez 12-20 dollars is not = 8

OpenStudy (mrnood):

@MissyMousey They are not instructions - they are possible answers Although which schoolkid doesn't know how much change he'll get for from $20 after buying four $3 items?

OpenStudy (anonymous):

20-12=8

OpenStudy (anonymous):

@MrNood Well, that makes more sense than directions! XD